# Efficient and Universally Accessible Cross-Chain Options without Upfront Holder Collateral

### v2

@@ IV-A1 Transfer Holder’s Position
Suppose Alice reaches an agreement with Carol to transfer the holder position on or before time $T_{H}$, with a charge of HolderFee on $Chain_{C}$. Carol performs $\text{KeyGen}(1^{\lambda})\rightarrow(pk_{C},sk_{C})$ to generate a new transfer key pair. Carol deposits HolderFee in $Contract_{C}$. This contract requires a signature of $m=(a,p)$, where message payload $p=(\text{Carol}.address,H(C),pk_{C})$, signed by $sk_{A}$ to unlock and transfer the HolderFee to Alice. $H(C)$ represents a new hashlock chosen by Carol, which is used to exercise the option if Carol obtains the holder position. Besides, $Contract_{C}$ records $H(A)$, specifying that HolderFee is refunded to Carol if Carol can reveal $A$ (meaning that Alice has exercised the option). Instead withdraw immediately, after Alice reveals a signature to redeem HolderFee, she must wait for $3\Delta$ to elapse. We refer to this period as the Withdrawal Delay Period. The protocol consists of two phases, Figure 1 illustrates the position transferring process:
@@
If all parties perform honestly, Alice is able to receive HolderFee and holder is changed to Carol. However, there are possible contingent events or dishonest scenarios:
@@
  • If different signatures with the same message address $\sigma_{m^{\prime}}\neq\sigma_{m}$, are submitted on $Chain_{A}$ and $Chain_{B}$ (e.g., if Alice submits two different signatures or sells the option to multiple parties), any one can call $\text{Extract}(pk,m^{\prime},\sigma_{m^{\prime}},m,\sigma_{m})\rightarrow sk_{A}$ to get $sk_{A}$. $sk_{A}$ is the secret transfer key of Alice. Whoever gets $sk_{A}$ means that Alice misbehaves. We can use this as an evidence for fair settlement of funds.
@@
Timeouts. The transfer contract must be created no later than $T_{H}-3\Delta$, and the reveal phase should be completed by $T_{H}-2\Delta$ to ensure that the option can be transferred to Carol at $T_{H}$. In the consistency phase, if any misbehavior occurs, it should be reported to the contract by $T_{H}+\Delta$. If Bob does not claim assets on $Chain_{A}$ and $Chain_{B}$ with $sk_{A}$, then it implies transfer complete. Overall, a total transfer time of $4\Delta$ is required. In other words, the transfer protocol must initiate no later than $T_{E}-4\Delta$. The unlocking condition for $Contract_{C}$ is summarized in Table II.
@@ IV-A2 How misbehaviors are handled securely in the protocol
Here we show how this protocol handles misbehaviors and protect each party’s interests by ensuring a fair payoff for honest parties in red dashed lines of Figure 1. A more rigorous analysis is shown in Section -A1. First, we consider each party acting maliciously on their own.
  • If Alice provides two different signatures to different buyers, as shown in , Bob can extract $sk_{A}$ and submit it to obtain $Asset_{A}$ and $Asset_{B}$, and Carol can reclaim the transfer fee with $sk_{A}$. In that case, Bob does not lose his $Asset_{B}$ and Carol does not lose her transfer fee.
  • If Alice reveals $A$ at the same time during the transfer process, as shown in , Carol can use $A$ to reclaim HolderFee. She does not lose anything. The option is exercised, and swap happens between Alice and Bob.
@@ IV-A3 Transfer Writer’s Position
Transferring the writer’s position is similar but simpler because Bob does not possess the preimage of the hashlock. Bob, with the transfer key pair $(pk_{B},sk_{B})$, can sign the message $m=(a,(\text{Dave}.address,pk_{D}))$ using $sk_{B}$ to collect the transfer fee, where $pk_{D}$ is a new transfer key for Dave. Transferring writer’s position does not update the hashlock used in the option exercise. Thus, Alice’s option is not influenced except the change of new option writer.
@@ IV-B Holder Collateral-Free Cross-Chain Options
Suppose Alice and Bob reach an agreement that Alice pays Bob $P$ as a premium on a chain denoted by $Chain_{P}$. The option takes effective at $T_{A}$ meaning that Alice obtains the right to exchange $Asset_{A}$ for $Asset_{B}$ before $T_{E}$. Bob’s guarantee is $Asset_{G}$.
The protocol involves two kinds of asset settlement: first for option establishment (or activation, we use them interchangably) and second for option exercise. We therefore introduce two secrets: (1) Activation secret $A$, used for Alice to pay the premium and activate the option; and (2) Exercise secret $B$, used for Alice to pay $Asset_{A}$ to Bob in exchange for $Asset_{B}$ when the option is exercised. Secret $A$ is generated by Alice and $B$ is generated by Bob.
@@
In the setup phase, Alice and Bob will establish this option similarly to a vanilla HTLC. Alice locks $P$ with a hashlock $H(A)$ in a contract on any chain. Bob creates two contracts, $Contract_{A}$ on $Chain_{A}$ and $Contract_{B}$ on $Chain_{B}$, which are used in the option. The option remains inactive until Alice reveals the activation secret $A$ before $T_{A}$, at which point the state updates to active and Bob gets Alice’s premium. $Contract_{A}$ holds Bob’s guarantee, $Asset_{G}$, until the option expires. If Alice exercises the option and Bob fulfills his obligation by revealing the exercise secret $B$, $Asset_{G}$ is refunded to Bob. If Bob fails to fulfills his obligation, $Asset_{G}$ will be transferred to Alice. $Contract_{B}$ holds Bob’s collateral, $Asset_{B}$, with hashlock $H(B)$.
@@
    Alice randomly selects a secret $A$ as activation secret, and computes its hash value $H(A)$. Bob generates $B$ and $H(B)$, which serve as the exercise secret and hashlock.
@@
    Alice locks $P$ with hashlock $H(A)$ on the agree-upon $Chain_{P}$ with timeout $T_{A}+\Delta$.
@@ IV-B1 Integration: Efficient Cross-Chain Options without Upfront Holder Collateral
Take Bob transferring writer to Dave as an example. It is similar to the Section IV-A1 with three notable differences. Suppose Bob reaches an agreement with Dave to transfer the writer position. Dave is able to buy Bob’s risky asset with its obligation at the price of WriterFee before or at $T_{W}$. Firstly, Dave should choose a new hashlock as the exercise secret. Similarly, Bob needs to use his private key $sk_{B}$ to sign Dave’s new hashlock $H(D)$, which means message $m=(a,(\text{Dave}.address,H(D),pk_{D}))$. Secondly, Alice can use Bob’s private key $sk_{B}$ to reclaim $Asset_{B}$ and guarantee, $Asset_{G}$. Thirdly, if Alice wants to exercise the option and makes the deposit after Bob reveals the signature during the transfer process, the transfer continues, and Dave should forward the signature to obtain the writer’s position. Dave should fulfill his obligation and reveal the exercise secret at $T_{W}-\Delta$ on $Contract_{A}$.
@@ IV-B2 Support Concurrent Bidding and Defending Against Phantom Bid Attacks
Since the support for concurrent bidding, our protocol can effectively defend against phantom bid attack. In phantom bid attack, an adversary creates multiple virtual buyers who offer higher prices but do not finalize the transfer. In the previous protocol [30] which attempts to transfer the option to a buyer sequentially, in face of such an attack, the option holder/writer cannot sell their positions in a reasonable time since the virtual buyers are exhausting the option transfer window.
@@ V SECURITY ANALYSIS
In this section, we analyze and prove the properties satisfied by our proposed protocol. Firstly, recall from the previous illustration that Alice/Carol is called the holder, and Bob/Dave is the writer, and Carol or Dave is required to pay a transfer fee to the corresponding seller. The payment $P$ is called the premium. $Asset_{G}$ deposited by Bob is called guarantee. transfer fee or any other asset escrowed in contracts are referred to as collateral.
@@ V-A Option Transfer Properties
  • If Alice is conforming and loses her position, then she will obtain either Carol’s collateral, Bob’s collateral and guarantee, or both.
@@
Evidently, Alice does not own the exercise secret, holder’s transfer is required to replace the holder address and transfer public key, and the inconsistency of two chains will not harm the interest of Bob. According to the Protocol IV-A, Bob cannot use the transfer private key of Alice, i.e. $sk_{A}$ to claim assets. Therefore, during the reveal phase and consistency phase, Bob is not required to participate and not allowed to make any change on $Contract_{A}$ and $Contract_{B}$. ∎
@@
By Lemma 1, Bob’s participation is not required during the holder transfer. If Alice and Carol are conforming, Carol will create $Contract_{C}$ contract and lock her collateral using signature of Alice before $T_{H}-3\Delta$. Alice will then reveal signature by $sk_{A}$ and call $reveal()$ on $Contract_{C}$ at $T_{H}-2\Delta$. An honest Carol will forward the signature, setting the holder to Carol. Alice can then wait for $3\Delta$ withdrawal delayed period to obtain the collateral, while the writers of $Contract_{A}$ and $Contract_{B}$ are still Bob, Bob maintains the writer’s position. During the process where Bob transfers his position to Dave, if both parties are conforming, Bob will not expose two different signatures. After $T_{W}+\Delta$, Bob will not be obstructed and will surely obtain Dave’s collateral. Meanwhile, Dave can submit $\sigma_{m}$ between $T_{W}-\Delta$ and $T_{W}$ to $Contract_{A}$ and $Contract_{B}$ to change the writer, Alice retaining the holder position. ∎
@@
Protocol IV-B2 satisfies transfer isolation: Alice and Bob can simultaneously and separately transfer their positions to Carol and Dave, respectively. This means that transferring holder and the transferring writer can proceed concurrently.
@@ V-B Option Properties
Protocol IV-B satisfies option correctness: If both the Alice and Bob are conforming, then if Alice does not exercise the right, Alice doesn’t lose the $Asset_{A}$ and Bob doesn’t lose the $Asset_{G}$ and $Asset_{B}$; or if Alice exercise the right, then Alice will receive $Asset_{B}$ and Bob will receive $Asset_{A}$ and $Asset_{G}$.
@@ VI EVALUATION
Implementation. We implemented, tested, and evaluated our proposed protocol, Efficient Cross-Chain Options without Upfront Holder Collateral, in Section IV-B1. We evaluated a cross-chain over two ERC-20 tokens on Ethereum Virtual Machine (EVM)-compatible chains. The experiment was conducted within the latest EVM and the Solidity Compiler version 0.8.22. The on-chain transaction cost (gas consumption) in this environment is identical to the current Ethereum network. DAPS can be implemented by algorithms like Schnorr, DSA, and EdDSA, which are natively supported by most blockchains. Implementing DAPS is straightforward, as signature algorithms inherent in blockchains can directly support it. We employ the same signature algorithm as Ethereum, utilizing the secp256k1 curve and the Elliptic Curve Digital Signature Algorithm (ECDSA) [35]. Since EVM does not support direct on-chain verification of a public-private key pair, we implement the proof of the private key $sk$ by signing a specific message with it. The corresponding codes are provided in Section -B.
@@
Expected Transfer Time Evaluation. We compared our work with that of Engel and Xue [30], since currently, their protocol is the only one that achieves transferability without the need for a trusted third-party cross-chain bridge. Assuming that the probability of an option being transferred and ultimately finalized within the current network is $p$, the total number of transfers is $X$. Then, in their protocol, $X$ follows a geometric distribution, i.e. $X\sim G(p)$. The relationship between the expected successful transfer time and the successful transfer probability of each phase is illustrated in Figure 3a and Figure 3b, where mutate lock phase in Engel and Xue’s protocol to corresponds to the reveal phase in our protocol. When a large number of malicious nodes exist in the current network, say, the finalization probability is 10%, the duration of the mutate lock phase and the consistency phase in their protocol becomes significantly prolonged, reaching $45\Delta$, which is approximately equivalent to 2 days in Bitcoin. By initiating the replace phase earlier and consolidating the mutate and consistency phases, we significantly reduce the duration of these phases.
@@
Gas Consumption Evaluation. Figure 4 lists the gas consumption for contract deployment, option operations, and gas consumption in different phases [^5] In the transfer failure case, we only consider the gas consumption of conforming parties., where gas price is $1$ Gwei (average price on 24 Feb, 2025) [36] and ETH price is \$$\$2313.17$ (24 Feb, 2025).
@@
Notably, as shown in Figure 4a, compared to Engel and Xue’s protocol, the gas consumption for the holder transfer process has significantly decreased from 714,867 to 510,857 gas (a reduction of approximately 28.5% for successful transfers). For failed transfers, the gas consumption decreases from 330,350 gas to 248,388 gas (a reduction of approximately 3.4%). The gas consumption of the transferring writer also decreases to a similar level. The gas consumption for exercising an option increases from 96,916 to 145,337, while the gas consumption for abandonment decreased. This is because, during an exercise, Alice needs to deposit funds and Bob must fulfill the request by revealing the exercise secret. In contrast, for an abandonment, Bob only needs to perform a refund operation.
@@
Figure 4b illustrates the gas consumption of a successful transfer across different phases. In our protocol, the reveal phase only requires the seller to reveal a signature in one contract, significantly reducing gas consumption compared to the mutate and replace/revert phases, lowering the gas for the holder and the writer to 123,158 and 123,435 gas, respectively. However, gas usage in the consistency phase is higher than that of their protocol, as we verify signatures in both contracts to ensure consistency.
The gas consumption for contract deployment in our protocol is generally higher compared to Engel and Xue’s protocol due to additional security measures and DAPS support. For instance, the deployment costs for $Contract_{A}$ and $Contract_{B}$ in our protocol are 2,549,610 and 2,220,156 gas, respectively.
@@
Nonetheless, the gas costs are acceptable, although some operations increase gas usage. Our protocol implements more secure operations and reduces the costs during transfers and abandonment. On the other hand, gas fees for each operation (usually less than \$6) are insignificant compared to the option’s value. Although gas costs for exercising options and deploying contracts are higher, this expense ensures efficient and robust transfers. Additionally, it also eliminates the holder’s collateral, which can be invested as other liquidity and increaset the capital efficiency. This addresses traders’ primary concerns.

### v4

@@ IV-A1 Transfer Holder’s Position
Suppose Alice reaches an agreement with Carol to transfer the holder position on or before time $T_{H}$, with a charge of HolderFee on $Chain_{C}$. Carol performs $\text{KeyGen}(1^{\lambda})\rightarrow(pk_{C},sk_{C})$ to generate a new transfer key pair. Carol deposits HolderFee in $Contract_{C}$. This contract requires a signature of $m=(a,p)$, where message payload $p=(\text{Carol}.address,H(C),pk_{C})$, signed by $sk_{A}$ to unlock and transfer the HolderFee to Alice. $H(C)$ represents a new hashlock chosen by Carol, which is used to exercise the option if Carol obtains the holder position. Besides, $Contract_{C}$ records $H(A)$, specifying that HolderFee is refunded to Carol if Carol can reveal $A$ (meaning that Alice has exercised the option). Instead of withdrawing immediately, after Alice reveals a signature to redeem HolderFee, she must wait for $3\Delta$ to elapse. We refer to this period as the Withdrawal Delay Period. The protocol consists of two phases, Figure 1 illustrates the position transferring process:
@@
If all parties perform honestly, Alice is able to receive HolderFee and the holder is changed to Carol. However, there are possible contingent events or dishonest scenarios:
@@
  • If different signatures with the same message address $\sigma_{m^{\prime}}\neq\sigma_{m}$, are submitted on $Chain_{A}$ and $Chain_{B}$ (e.g., if Alice submits two different signatures or sells the option to multiple parties), anyone can call $\text{Extract}(pk,m^{\prime},\sigma_{m^{\prime}},m,\sigma_{m})\rightarrow sk_{A}$ to get $sk_{A}$. $sk_{A}$ is the secret transfer key of Alice. Whoever gets $sk_{A}$ means that Alice misbehaves. We can use this as evidence for fair settlement of funds.
@@
Timeouts. The transfer contract must be created no later than $T_{H}-3\Delta$, and the reveal phase should be completed by $T_{H}-2\Delta$ to ensure that the option can be transferred to Carol at $T_{H}$. In the consistency phase, if any misbehavior occurs, it should be reported to the contract by $T_{H}+\Delta$. If Bob does not claim assets on $Chain_{A}$ and $Chain_{B}$ with $sk_{A}$, then it implies transfer is complete. Overall, a total transfer time of $4\Delta$ is required. In other words, the transfer protocol must initiate no later than $T_{E}-4\Delta$. The unlocking condition for $Contract_{C}$ is summarized in Table II.
@@ IV-A2 How misbehaviors are handled securely in the protocol
Here we show how this protocol handles misbehaviors and protects each party’s interests by ensuring a fair payoff for honest parties in red dashed lines of Figure 1. A more rigorous analysis is shown in the full version [peng_full_version] of this paper. First, we consider each party acting maliciously on its own.
  • If Alice provides two different signatures to different buyers, as shown in , Bob can extract $sk_{A}$ and submit it to obtain $Asset_{A}$ and $Asset_{B}$, and Carol can reclaim the transfer fee with $sk_{A}$. In that case, Bob does not lose his $Asset_{B}$, and Carol does not lose her transfer fee.
  • If Alice reveals $A$ at the same time during the transfer process, as shown in , Carol can use $A$ to reclaim HolderFee. She does not lose anything. The option is exercised, and a swap happens between Alice and Bob.
@@ IV-A3 Transfer Writer’s Position
Transferring the writer’s position is similar but simpler because Bob does not possess the preimage of the hashlock. Bob, with the transfer key pair $(pk_{B},sk_{B})$, can sign the message $m=(a,(\text{Dave}.address,pk_{D}))$ using $sk_{B}$ to collect the transfer fee, where $pk_{D}$ is a new transfer key for Dave. Transferring the writer’s position does not update the hashlock used in the option exercise. Thus, Alice’s option is not influenced except for the change of the new option writer.
@@ IV-B Holder Collateral-Free Cross-Chain Options
Suppose Alice and Bob reach an agreement that Alice pays Bob $P$ as a premium on a chain denoted by $Chain_{P}$. The option takes effect at $T_{A}$ meaning that Alice obtains the right to exchange $Asset_{A}$ for $Asset_{B}$ before $T_{E}$. Bob’s guarantee is $Asset_{G}$. The protocol involves two kinds of asset settlement: first for option establishment (or activation, we use them interchangeably) and second for option exercise. We therefore introduce two secrets: (1) Activation secret $A$, used for Alice to pay the premium and activate the option; and (2) Exercise secret $B$, used for Alice to pay $Asset_{A}$ to Bob in exchange for $Asset_{B}$ when the option is exercised. Secret $A$ is generated by Alice and $B$ is generated by Bob.
@@
In the setup phase, Alice and Bob will establish this option similarly to a vanilla HTLC. Alice locks $P$ with a hashlock $H(A)$ in a contract on any chain. Bob creates two contracts, $Contract_{A}$ on $Chain_{A}$ and $Contract_{B}$ on $Chain_{B}$, which are used in the option. The option remains inactive until Alice reveals the activation secret $A$ before $T_{A}$, at which point the state updates to active and Bob gets Alice’s premium. $Contract_{A}$ holds Bob’s guarantee, $Asset_{G}$, until the option expires. If Alice exercises the option and Bob fulfills his obligation by revealing the exercise secret $B$, $Asset_{G}$ is refunded to Bob. If Bob fails to fulfill his obligation, $Asset_{G}$ will be transferred to Alice. $Contract_{B}$ holds Bob’s collateral, $Asset_{B}$, with hashlock $H(B)$.
@@
    Alice randomly selects a secret $A$ as the activation secret, and computes its hash value $H(A)$. Bob generates $B$ and $H(B)$, which serve as the exercise secret and hashlock.
@@
    Alice locks $P$ with hashlock $H(A)$ on the agreed-upon $Chain_{P}$ with a timeout $T_{A}+\Delta$.
@@ IV-B1 Integration: Efficient Cross-Chain Options without Upfront Holder Collateral
Take Bob transferring the writer’s position to Dave as an example. It is similar to the Section IV-A1 with three notable differences. Suppose Bob reaches an agreement with Dave to transfer the writer position. Dave is able to buy Bob’s risky asset with its obligation at the price of WriterFee before or at $T_{W}$. Firstly, Dave should choose a new hashlock as the exercise secret. Similarly, Bob needs to use his private key $sk_{B}$ to sign Dave’s new hashlock $H(D)$, which means message $m=(a,(\text{Dave}.address,H(D),pk_{D}))$. Secondly, Alice can use Bob’s private key $sk_{B}$ to reclaim $Asset_{B}$ and guarantee, $Asset_{G}$. Thirdly, if Alice wants to exercise the option and makes the deposit after Bob reveals the signature during the transfer process, the transfer continues, and Dave should forward the signature to obtain the writer’s position. Dave should fulfill his obligation and reveal the exercise secret at $T_{W}-\Delta$ on $Contract_{A}$.
@@ IV-B2 Support Concurrent Bidding and Defending Against Phantom Bid Attacks
Since the support for concurrent bidding, our protocol can effectively defend against phantom bid attack. In the phantom bid attack, an adversary creates multiple virtual buyers who offer higher prices but do not finalize the transfer. In TCO [30], which attempts to transfer the option to a buyer sequentially, in the face of such an attack, the option holder/writer cannot sell their positions in a reasonable time since the virtual buyers are exhausting the option transfer window.
@@ V SECURITY ANALYSIS
In this section, we analyze and prove the properties satisfied by our proposed protocol. Firstly, recall from the previous illustration that Alice/Carol is called the holder, and Bob/Dave is the writer, and Carol or Dave is required to pay a transfer fee to the corresponding seller. The payment $P$ is called the premium. $Asset_{G}$ deposited by Bob is called guarantee. transfer fee or any other asset escrowed in contracts is referred to as collateral.
@@ V-A Option Transfer Properties
  • If Alice is conforming and loses her position, then she will obtain either Carol’s collateral, Bob’s collateral, and guarantee, or both.
@@
Evidently, Alice does not own the exercise secret, holder’s transfer is required to replace the holder address and transfer public key, and the inconsistency of two chains will not harm the interest of Bob. According to the Protocol IV-A, Bob cannot use the transfer private key of Alice, i.e. $sk_{A}$ to claim assets. Therefore, during the reveal phase and consistency phase, Bob is not required to participate and is not allowed to make any changes on $Contract_{A}$ and $Contract_{B}$. ∎
@@
By Lemma 1, Bob’s participation is not required during the holder transfer. If Alice and Carol are conforming, Carol creates $Contract_{C}$ contract and lock her collateral using the signature of Alice before $T_{H}-3\Delta$. Alice will then reveal the signature by $sk_{A}$ and call $reveal()$ on $Contract_{C}$ at $T_{H}-2\Delta$. An honest Carol will forward the signature, setting the holder to Carol. Alice can then wait for $3\Delta$ withdrawal delayed period to obtain the collateral, while the writer of $Contract_{A}$ and $Contract_{B}$ is still Bob, Bob maintains the writer’s position. During the process where Bob transfers his position to Dave, if both parties are conforming, Bob will not expose two different signatures. After $T_{W}+\Delta$, Bob will not be obstructed and will obtain Dave’s collateral. Meanwhile, Dave can submit $\sigma_{m}$ between $T_{W}-\Delta$ and $T_{W}$ to $Contract_{A}$ and $Contract_{B}$ to change the writer, Alice retains the holder position. ∎
@@
Protocol IV-B2 satisfies transfer isolation: Alice and Bob can simultaneously and separately transfer their positions to Carol and Dave, respectively. This means that the transferring holder and the transferring writer can proceed concurrently.
@@ V-B Option Properties
Protocol IV-B satisfies option correctness: If both Alice and Bob are conforming, then if Alice does not exercise the right, Alice doesn’t lose $Asset_{A}$ and Bob doesn’t lose $Asset_{G}$ and $Asset_{B}$; or if Alice exercises the right, then Alice will receive $Asset_{B}$ and Bob will receive $Asset_{A}$ and $Asset_{G}$.
@@ VI EVALUATION
Implementation. We implemented, tested, and evaluated our proposed protocol [^5] Code is available in https://github.com/sfofgalaxy/cross-chain-options, Efficient Cross-Chain Options without Upfront Holder Collateral in Section IV-B1. We evaluated a cross-chain over two ERC-20 tokens on Ethereum Virtual Machine (EVM)-compatible chains. The experiment was conducted within the latest EVM and the Solidity Compiler version 0.8.22. The on-chain transaction cost (gas consumption) in this environment is identical to the current Ethereum network. DAPS can be implemented by algorithms like Schnorr, DSA, and EdDSA, which are natively supported by most blockchains. Implementing DAPS is straightforward, as signature algorithms inherent in blockchains can directly support it. We employ the same signature algorithm as Ethereum, utilizing the secp256k1 curve and the Elliptic Curve Digital Signature Algorithm (ECDSA) [35]. Since EVM does not support direct on-chain verification of a public-private key pair, we implement the proof of the private key $sk$ by signing a specific message with it.
@@
Expected Transfer Time Evaluation. We compared our work with TCO [30], since currently, only their protocol achieves transferability without the need for a trusted third-party cross-chain bridge. Assuming that the probability of an option being transferred and ultimately finalized within the current network is $p$, the total number of transfers is $X$. Then, in their protocol, $X$ follows a geometric distribution, i.e., $X\sim G(p)$. The expected successful transfer time and the transfer probability of each phase are illustrated in Figure 3a and Figure 3b. When many malicious nodes exist in the current network, say, the finalization probability is 10%, the duration of the mutate lock phase and the consistency phase in their protocol becomes significantly prolonged, reaching $45\Delta$, which is approximately equivalent to 2 days in Bitcoin. By initiating the replace phase earlier and consolidating the mutate and consistency phases, we significantly reduce the duration of these phases.
@@
Gas Consumption Evaluation. Figure 4 lists the gas consumption for contract deployment, option operations, and gas consumption in different phases [^6] In the transfer failure case, we only consider the gas consumption of conforming parties., where gas price is $1$ Gwei (average price) [36] and ETH price is \$$\$2313.17$ on 24 Feb, 2025.
@@
Notably, as shown in Figure 4a, compared to TCO, the gas consumption for the holder transfer process has significantly decreased from 714,867 to 510,857 gas (a reduction of approximately 28.5% for successful transfers). For failed transfers, the gas consumption decreases from 330,350 gas to 248,388 gas (a reduction of approximately 3.4%). The gas consumption of the transferring writer also decreases to a similar level. The gas consumption for exercising an option increases from 96,916 to 145,337, while the gas consumption for abandonment decreases. This is because, during an exercise, Alice needs to deposit funds, and Bob must fulfill the request by revealing the exercise secret.
@@
In contrast, for an abandonment, Bob only needs to perform a refund operation. Figure 4b illustrates the gas consumption of a successful transfer across different phases. In our protocol, the reveal phase only requires the seller to reveal a signature in one contract, significantly reducing gas consumption compared to the mutate and replace/revert phases, lowering the gas for the holder and the writer to 123,158 and 123,435 gas, respectively. However, gas usage in the consistency phase is higher than that of their protocol, as we verify signatures in both contracts to ensure consistency. The gas consumption for contract deployment in our protocol is generally higher compared to TCO due to additional security measures and DAPS support. The deployment costs for $Contract_{A}$ and $Contract_{B}$ in our protocol are about 2.5M and 2.2M gas.
@@
Nonetheless, the gas costs are acceptable, although some operations increase gas usage. Our protocol implements more secure operations and reduces the costs during transfers and abandonment. On the other hand, gas fees for each operation (usually less than \$6) are insignificant compared to the option’s value. Although gas costs for exercising options and deploying contracts are higher, this expense ensures efficient and robust transfers. Additionally, it also eliminates the holder’s collateral, which can be invested as other liquidity and increase the capital efficiency. This addresses traders’ primary concerns.